can you help me solve substitution y=3x and y-x=4
Answer by jim_thompson5910(35256)

Hint:

y = 3x, so you can replace the 'y' in y - x = 4 with 3x to get

y - x = 4

3x - x = 4

2x = 4

I'll let you finish in solving for x. Once you have the solution for x, use it to find the solution for y.